OpenStudy (anonymous):

(3/x-2) - (2/x+4) = 1

OpenStudy (anonymous):

do you mean \[\frac{ 2 }{ x-3 } - \frac{ 2 }{ x+4 } = 1\]

OpenStudy (phoenixfire):

\[{3 \over {x - 2}}-{2 \over {x+4}}=1\] Right?

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

PhoenixFire, yes! My computer's being super slow.

OpenStudy (anonymous):

Okay find the Common denominator? (x-2)(x+4)

OpenStudy (anonymous):

and multiply both sides by it

OpenStudy (anonymous):

Yep! I tried multiplying everything by (x-2) and (x+4).

OpenStudy (anonymous):

\[(x-2)(x+4) \left[ \frac{ 3 }{ x-2 } - \frac{ 2 }{ x+4 } =1\right]\]

OpenStudy (anonymous):

getting 3(x+4) - 2(x-2) = (x-2)(x+4)

OpenStudy (anonymous):

That's exactly what I got!

OpenStudy (anonymous):

3x+12-2x-4= Remember to foil this side

OpenStudy (anonymous):

Got that!

OpenStudy (phoenixfire):

it's 3x+12-2x+4... not -4

OpenStudy (anonymous):

3x+12-2x-4= x`2 +2x-8

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

When I figured it all out, I ended up having to use the quadratic formula. I got +4, too.

OpenStudy (anonymous):

it's x + 16 = x`2 +2x-8

OpenStudy (anonymous):

Got that!!

OpenStudy (anonymous):

get everything on one side

OpenStudy (anonymous):

x`2+x-24 = 0

OpenStudy (anonymous):

Then, it's 0=x^2+x-24. You can't factor that, can you?

OpenStudy (anonymous):

hmm it not easily factor-able so use quadratic formula

OpenStudy (anonymous):

But, it's not a real solution, is it?

OpenStudy (anonymous):

we can tell with the discriminate, it might be irrational number but not a nice clean answer

OpenStudy (anonymous):

\[\sqrt{1^{2}-4(1)(-24)}\]

OpenStudy (anonymous):

The question was to find all real values for x. I didn't think this was a real answer:\[\frac{ -1\pm \sqrt{97} }{ 2 }\]

OpenStudy (anonymous):

when you say REAL do you mean Real Numbers, because irrationals are real numbers

OpenStudy (anonymous):

True... I suppose I'll put that as my answer!

OpenStudy (anonymous):

it would not be real if the radical was negative

OpenStudy (anonymous):

I would

OpenStudy (anonymous):

unless it was a approx answer like a decimal

OpenStudy (anonymous):

Thank you! I'm glad I did it right! Lol.

OpenStudy (anonymous):

good job, Its nice to see students doing math well

OpenStudy (anonymous):

Yeah, I love whenever people help me and I help them with math. It's kinda fun!

OpenStudy (anonymous):

I think so too

OpenStudy (anonymous):

Thanks for the help! :)

OpenStudy (anonymous):

NP